3. A force of F(x)=x²-cos(3x) +2, x is in meters, acts on an object. What is the work required to move the object from x=3 to x=7?

1
Expert's answer
2022-05-26T17:05:17-0400
"W=\displaystyle\int_{3}^{7}(x^2-\cos(3x) +2)dx"

"=[\dfrac{x^3}{3}-\dfrac{1}{3}\sin (3x)+2x]\begin{matrix}
 7 \\
 3
\end{matrix}"

"=\dfrac{343}{3}-\dfrac{1}{3}\sin (21)+14"

"-(\dfrac{27}{3}-\dfrac{1}{3}\sin (9)+6)"

"=\dfrac{340}{3}-\dfrac{1}{3}\sin (21)+\dfrac{1}{3}\sin (9)"